Your company has arranged a revolving credit agreement for up to $78 million at an interest rate of 1.47 percent per quarter. The agreement also requires your company to maintain a compensating balance of 4 percent of the unused portion of the credit line, to be deposited in a non-interest bearing account. Your company's short-term investment account at the same bank pays an interest rate of 0.61 per quarter. What is the effective annual interest rate if your company borrows $43 million for one year? Please show work!!!

Answers: 6.70%, 6.01%, 6.09%, 6.56%, 7.66%
Answer #1

Total Credit = $78,000,000

Used Credit = $43,000,000

Unused Credit = $78,000,000 - $43,000,000 = $44,000,000

Compensating Balance = 4% of unused credit = 4% * $44,000,000 = $1,400,000.00

Quarterly Cost of Used Credit = 1.47% * $43,000,000 = $632,100.00

Quarterly Cost of compensating Balance (Opportunity Cost) = 0.61% * $1,400,000 = $8,540.00

Total Cost of borrowing = $632,100 + $8,540 = $640,640.00

Effective quarterly interest = 640640/43000000 = 1.4899%

EAR = (1+1.4899%)^4 – 1 = 0.0609 or 6.09%

Effective annual interest rate is 6.09%